# The Multi-Query Paradox in Zeroth-Order Optimization

### v1

@@ A.1 Proofs
A.2 Experiment Set Up
Objective Functions
For the strongly convex quadratic function, $f(x)=\frac{1}{2}x^{T}Ax+b^{T}x$, the positive definite matrix $A$ was generated by first creating a random matrix $M$ with entries sampled from a standard normal distribution. $A$ was then computed as $A=M^{T}M+\epsilon I$, where $\epsilon$ is a small positive constant to ensure the matrix is well-conditioned. The vector $b$ was also sampled from a standard normal distribution.
For the convex logistic regression task, synthetic data was generated by sampling feature vectors $a_{i}\in\mathbb{R}^{1000}$ from a standard normal distribution. A ground-truth weight vector $w_{true}$ was randomly created, and the binary labels $y_{i}\in\{-1,1\}$ were then generated based on the sign of $a_{i}^{T}w_{true}$.
The non-convex objective was the standard Rosenbrock function.
The stochastic convex objective used a logistic regression loss, where mini-batches were sampled at each iteration to compute the function value.
Implementation and Evaluation
We evaluated the ZO-Avg and ZO-Align estimators using various query block sizes $q$. Step sizes $\eta_{t}$ were set according to the optimal values derived in our theory: for ZO-Avg, $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$, and for ZO-Align, $\eta_{t}=1/L$. For the stochastic setting, a standard diminishing step-size rule was applied. Each algorithm was tested under two total query budgets: a large budget of K=20,000 and a constrained budget of K=500. All reported results are an average of 10 independent runs to ensure reliability.

### v3

@@ B.2 Proof of Proposition LABEL:prop:biasedness_zo_align
Appendix C Experiment Setup
Classical objectives.
All four classical problems use dimension $d=1000$.
For the strongly convex quadratic, $f(x)=\tfrac{1}{2}x^{\top}Ax+b^{\top}x$, we first sample a matrix $M\in\mathbb{R}^{d\times d}$ with i.i.d. $\mathcal{N}(0,1)$ entries and set $A=M^{\top}M+\epsilon I_{d}$ with a small $\epsilon>0$ to ensure good conditioning. The vector $b\in\mathbb{R}^{d}$ is sampled from $\mathcal{N}(0,I_{d})$.
For the convex logistic regression objective, we generate synthetic data by sampling feature vectors $a_{i}\in\mathbb{R}^{d}$ i.i.d. from $\mathcal{N}(0,I_{d})$. We draw a ground-truth weight vector $w_{\mathrm{true}}\sim\mathcal{N}(0,I_{d})$ and assign binary labels $y_{i}\in\{-1,1\}$ according to $\mathrm{sign}(a_{i}^{\top}w_{\mathrm{true}})$ (with ties broken arbitrarily). The objective is the empirical logistic loss $\tfrac{1}{m}\sum_{i=1}^{m}\log(1+\exp(-y_{i}a_{i}^{\top}x))$.
The non-convex objective is the standard Rosenbrock function $f(x)=\sum_{i=1}^{d-1}\bigl[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}\bigr]$.
For the stochastic convex objective, we consider an $\ell_{2}$-regularized logistic regression loss,
$f(x)=\mathbb{E}_{(\mathbf{w},y)\sim\mathcal{D}}\bigl[\log(1+\exp(-y\,\mathbf{w}^{\top}x))\bigr]+\tfrac{\rho}{2}\|x\|^{2},$
where $\mathcal{D}$ is a synthetic distribution generated in the same way as for the convex logistic regression. In the algorithm, this expectation is approximated using mini-batches of examples resampled at each iteration.
Classical experiments: implementation details.
For each classical objective we run both ZO-Avg and ZO-Align with several query block sizes $q$ and two total query budgets $K\in\{500,20{,}000\}$, as described in Section 6. Unless otherwise stated, we use the idealized one-sided finite-difference estimator with a fixed smoothing parameter $\mu=10^{-6}$. For ZO-Avg and ZO-Align in the deterministic settings, the step sizes are set according to the theoretically optimal choices in Proposition 4.1, namely $\eta_{t}=\tfrac{q_{t}}{L(q_{t}+d+1)}$ for ZO-Avg and $\eta_{t}=1/L$ for ZO-Align, where $L$ is the Lipschitz constant of the gradient (for the quadratic objective, $L$ equals the largest eigenvalue of $A$). In the stochastic convex case we use diminishing step sizes of the form $\eta_{t}=\eta_{0}/\sqrt{t+1}$, with $\eta_{0}$ tuned by a small grid search. All reported curves are averaged over 10 independent runs with different random seeds.
High-dimensional language model fine-tuning.
For the high-dimensional experiments we fine-tune the Qwen3-0.6B model on the SST-2 and CB sentence classification datasets using zeroth-order optimization on the cross-entropy loss. We follow a standard sequence-classification protocol: each input sentence (or sentence pair) is fed to the model, the representation of the special classification token is passed through a linear head to produce class logits, and the cross-entropy with respect to the ground-truth labels is used as the objective. All trainable parameters (including the classification head) are flattened into a single vector in $\mathbb{R}^{d}$, where $d$ is on the order of the total number of parameters of the fine-tuned model.
At each zeroth-order update step we sample $q$ i.i.d. Gaussian directions and form the direction matrix $U\in\mathbb{R}^{d\times q}$. ZO-Avg uses the standard averaging estimator based on these $q$ directions. For ZO-Align we use the diagonal approximation described in Section 6.2: instead of explicitly forming and inverting the full Gram matrix $U^{\top}U$, we compute only its diagonal $D=\mathrm{diag}(U^{\top}U)$ and apply the update $\hat{g}_{\mathrm{ALG\text{-}diag}}(x)=UD^{-1}v$, where $v\in\mathbb{R}^{q}$ collects the one-sided finite-difference directional derivatives. This approximation is motivated by the high-dimensional analysis in Section 5, which shows that $U^{\top}U$ concentrates near a scaled identity when $d\gg q$.
We fix a total query budget of K=20, [^0] K=20{,}000 function evaluations and consider query block sizes $q\in\{10,15,50,100,500\}$. For each estimator (ZO-Avg and diagonal ZO-Align) and each choice of $q$, we run the optimizer until the budget $K$ is exhausted. A constant step size is used within each run; for fairness, we tune this step size separately for each (estimator, $q$) pair by a small logarithmic grid search.